OpenStudy (anonymous):

I'll fan and give a medal! :) Which conjunction or disjunction is equivalent to the open sentence 4 – 2|n + 6| ≥ 2? A. n > –5 or n < –7 B. n + 6 ≤ 1 and n + 6 ≥ –1 C. –5 < n or n < –7 D. n + 6 ≤ –1 and n + 6 ≥ 1

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (anonymous):

@amoodarya @KirbyLegs

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf 4 - 2|n + 6| \ge 2\implies - 2|n + 6|\ge 2+4 \implies - 2|n + 6|\ge -2\\ \quad \\ \textit{notice, we'll multiply by a negative value,thus } \color{red}{\textit{flipping the sign}}\\ \quad \\ - 2|n + 6|\ge -2\implies |n+6|\le \cfrac{-2}{-2}\implies |n+6|\le 1\\ \quad \\ |n+6|\le 1\implies \begin{cases} +(n+6)\le 1\implies n+6\le 1\\ \quad \\ \bf -(n+6)\le 1\implies n+6\ge -1 \end{cases} \)

OpenStudy (anonymous):

So the answer is D, correct?

OpenStudy (anonymous):

@jdoe0001